# On the Robustness of Machine Learning Models in Predicting Thermodynamic Properties: a Case of Searching for New Quasicrystal Approximants

### v1

@@ 1 Introduction
Over the last decade, artificial intelligence (AI) has become a powerful tool for various tasks in numerous fields, such as natural language processing (NLP) Roser (2022); Zhou et al. (2023), computer vision (CV) Roser (2022); Haenlein and Kaplan (2019), natural sciences Han et al. (2023); Ivanenkov et al. (2023), etc. With the advent of neural networks, which are, unfortunately, uninterpretable black boxes Alwosheel et al. (2021), the scientific and industrial communities began to pay attention to the issues of stability and robustness Hamon et al. (2020); Kejriwal et al. (2024). First, let us define what we call ’robustness’ hereafter. It is known, that negligible changes in the input data (both for training model or inference model) can provoke uncontrolled changes in the output of AI. Since the AI-assisted solutions affect real world decisions, this output, in turn, can lead to erroneous, costly, unethical, or even dangerous results. Thus, the robustness of machine learning (ML) algorithms characterize absense of the aforementioned situations and seems to be one of the key issues in the design of AI-assisted systems. The similar problems were discussed in detail by researchers in the NLP Goyal et al. (2023); Wang et al. (2021) and CV Liu and Jin (2023); Li et al. (2024) domains, but it can certainly arise wherever neural networks are used, including materials science.
@@
Indeed, AI fever has not spared materials science either: ML has become a solid part of the pipeline for searching for new materials by high-throughput thermodynamic stability assessments within the vast chemical search spaces. The most recent approaches in the field are resting upon testing chemical modifications of existing materials through their compositional/chemical modifications Merchant et al. (2023); Chen et al. (2024) instead of generation of crystal structures from scratch. Despite the fact that many modern computational datasets, such as that of Materials Project Jain et al. (2013) and AFLOWLib Curtarolo et al. (2012a), can be used to pre-train AI models for their subsequent applications, fine-tuning stage or even more sophisticated routines (e.g. active learning Merchant et al. (2023); Yuan et al. (2023)) are generally required for disordered materials. In such approaches, the more complex crystal structure, the more complex the target search space. For defects in 2D materials modelled within the supercell approach Huang et al. (2023), the authors provided combinations of structured composition/configuration spaces at low defect contents and random structures subsamples at higher ones. Nevertheless, it was recently demonstrated Eremin et al. (2024) that physicochemical intuition can provide an acceptable quality of model predictions at low computational costs. Thus, training data sampling not only can play a major role in most effectively combining exhaustive and high-precision (such as density functional theory) calculations and fast screening data-driven approaches, but is also a point for potential instability: small errors and inaccuracies in the composition of the training dataset can lead to unnecessary and unsuccessful resource-intensive ab-initio computational and expensive real-life experiments.
@@
In our recent research Eremin et al. (2022), we addressed thermodynamic stability assessments for the disordered Sc-rich complex (ca. 140 atoms/cell) intermetallics by combining density functional theory calculations and data-driven approaches. During the data collection, there was obtained considerable differences in thermodynamic favorability of defects of certain types. Despite this fact, all the data collected was used for model training. This allowed graph neural networks (GNN) and descriptor-based ML models to gain sufficient generalization ability. In turn, at the inference stage, the complete composition/configuration space was built only for the thermodynamically favored defects resulting in an appropriate complexity of target search space. Thus, the peculiarity of the developed approach was that structures with disorder of different types were used at the training/validation and inference stages.
@@
The aim of the current contribution is to estimate the unstability of screening pipeline and evaluate robustness of different backbones of it. The main goal is evaluating impact of the less relevant to the target domain data impact on the results. Using the data and the approach presented in Eremin et al. (2022), we have genenerated 13 subdatasets based on the disordering level and thus have modeled the changes in the training data. It is shown that the Random Forest algorithm Breiman (2001) is robust and stable against the training data changes, while the Allegro Musaelian et al. (2023) neural network is not. Its predictions could became completely different; the assessment was done not only in a quantitative root squared error (RMSE) way, but as well in a qualitative manner. We have also demonstrated that the pre-training on the corresponding Aflow dataset Curtarolo et al. (2012a) slice increases stability and introduced an elementary sequential training trick to increase algorithm robustness without involving additional data.
@@ 3.1 Overall description
The dataset consists of doped Sc-M intermetallic systems, scandium-rich Mackay-type quasicrystal approximants . The dopants M are four metals: Pt, Pd, Ir and Rh, and thus four corresponding subdatasets were formed. On the A4 and A5 sites corresponding to the intercluster regions, the original Sc60M13 crystal structure was disordered by adding vacancies or Sc substitutions by metal M with a step of 2 atoms. Then the structures with low Sc content (less than 0.667 for Pt and Pd, less than 0.733 for Ir and Rh) were excluded from the furter consideration due to the well-known competing phases. For th erest part of the search space, one random realization of crystal structure with each of the remaining compositions was chosen and computed using Vienna Ab initio Simulation Package (VASP) Kresse and Hafner (1993) Density Function Theory (DFT)-solver.
@@ 3.2 The term "limit"
The DFT-calculations showed that defects on the A4 site are more energetically favorable than that on A5 one, and thus we used numbers of defects on A5 as a quantitatively measure of sample similarity. So, we introduce the ’limit’ term corresponding to the maximum number of defects on the A5 site allowed to add to a certain subdataset. The maximum limit in the dataset equals multiplicity of the A5 site, 24, and the minimum is 0. Thus, the entire dataset can be represented as 13 nested datasets with the limit from 0 to 24. The obtained subdataset sizes depending on the limit are shown in Figure 1.
@@ 3.3 Splits
Secondly, each of the four Sc-Pd, Sc-Pt, Sc-Ir, and Sc-Rh datasets is divided into train, validation, and test samples in the same manner: the test set consists only of limit=0 part, while train and val contain all the other systems with the number of defects on A5 from 0 to 24. Train/val partition was made once for the full (limit=24) dataset and than applied accordingly to the subdatasets with limits less that 24. The precise sizes of all the data samples obtained are represented in Table 1
@@ 3.4 Target energies
It is worth noting that the energy above the hull is, by design, positive for most systems, and that the ultimate goal of the screening pipeline is to search for crystals that have negative (or at least slightly positive) energy: only such systems meets synthesisability restrictions. Figure 2 illustrates the distribution of the energies above the hull in the Sc-Pd combined dataset.
Figure 2: Histogram of energis above the hull histogram for the Sc-Pd dataset
@@ 4 Methods
For each architecture and for each dataset, we used the following procedure: 13 models was trained (one for each limit=$x$ subdataset) and compared on the single test sample. Than an additional comparison was made on the inference dataset
@@ 4.3 Training
For classical models, 5-fold cross-validation was used the predictions of which were then averaged, as for the neural network we used a described in 3.3 piece of data for validation purposes. Root mean square error (RMSE) 3 was used as error function for training and validation steps. Random seeds were fixed for all experiments.
@@ 5.1 Test
Even though RMSEs of all classical Gradient Boosting runs (see Figure 3) does not significantly vary (XGBoost) or have a clearly visible pattern (CatBoost and LightGBM)
@@
, all of the chosen Gradient Boosting algorithms perform worse than Random Forest approach, which can be seen from Figure 4
@@
Random Forest have much lower test RMSE compared to that of the tested Gradient Boosting models. Random Forest shows test RMSE of ca. 0.01 eV/atom and Gradient Boosting models’ scores are of ca 0.02 eV/atom. The latter have a tendency to overfit on the subdatasets with bigger limits (while test sample corresponds to limit=0) and their predictions are highly biased. Moreover, we checked made and checked the predictions of the trained Gradient Boosting models for the inference set. The results obtained were rather unrealistic and inconsistent. For this reason, Gradient Boostings were excluded from further consideration and only the Random Forest model is used.
Random Forest provides stable results, i.e. the test RMSEs are found to fluctuate within a few meV/atom for three out of four datasets (except for the Sc-Pd systems) as shown in Figure 5.
@@
In Figure 6, the SimilarityScores matrices also illustrates the Random Forest stability: sharp changes are common only for the Sc-Pt systems.
@@
Different neural network training approaches perform about the same RMSE ranges of 0.010 - 0.015 eV/atom on the test set. As it is shown in Figure 7, the self-pre-trained Allegro model’s errors are always smaller then that of the ordinarily trained one. The Aflow-pretrained version, in turn, shows bigger errors.
@@
In contrast to the similar RMSE values of the models tested, the Similarity Score shows pronounce differences between the model predictions at a qualitative level. The Aflow- (Figure 8) and self-pretrained (Figure 9) Allegro models provide heatmaps that illustrates larger Similarity Scores then that on the ordinarily trained network (Figure 10).
@@
Moreover, the light areas are much more clearly located. If a light bar appears after a certain limit, it means that it is likely to continue at the next limits. For some metals, there is a pattern where the model gets good results up to a certain limit comprising the relevant data only. Then, the model shows a worse result due to getting a little less relevant data. However, the model being provided by a large amount of both the relevant (limit=0,) and less relevant (limit=$x$, $x>0$) data shows better results again. It can be assumed that by perceiving the overall data it increases the robustness of predictions.
@@ 5.2 Inference
The comparison between the limit=24 models and all the others shows that the Random Forest is robust and the Allegro model is not. The recall metric computed considering the limit=24 predictions as the ground truth is always 1 for Random Forests (see Figure 11 a) and vary from 0 to 1 for Allegro (see Figure 11 b).
@@
The pairwise comparison of limits described in 4.5 highlights apparent difference between the models, trained on different subdatasets with limit=$x$ Random Forest behaves stable, showing not the chaotic changes, but a visible transition (see Figure 12) from different to similar experiments. The empty graphs and solid-colored heatmaps show that there are no structures predicted below the convex hull.
@@
The heatmaps indicate the instability of neural networks, as this behavior is not observed in the results of the Random Forest. Despite the fact that the results may vary, they are not random, and there is some kind of dependence factor. As it can be seen in Figure 13, there is a trade-off between the relevance of certain datapoints and their ammounts provided to the model. In the middle range of the introduced limits, a small change in the limit leads to large changes in the results. This observation clearly illustrate the folloing principle – training data should be either small and strictly relevant or big and less relevant. From the dependencies obtained for Random Forest, this can also be seen.
@@
As shown in Figure 14, the trade-off is definitely more visible and after a certain limit on each metal there is a great similarity between structures that are predicted below the convex hull. This also shows that there is a certain “relevance threshold” for palladium and platinum. With pretrained network, we can overcome unstable behavior of neural-network-based approaches without generation of additional data. Thresholds differ from those seen in the pictures of sequential training and regular training. This is most likely due to the fact that the model was already familiar with similar data patterns, and its ordinar training requires less data particular for a certain task. For rhodium, the threshold is also apparently observed, which shows that when training on data with a limit of 6 or more, the model predicts structures below the hull. The results for iridium do not allow us to unambiguously assess the effectiveness of the method proposed.

### v2

@@ 1 Introduction
Over the last decade, artificial intelligence (AI) has become a powerful tool for various tasks in numerous fields, such as natural language processing (NLP) Roser (2022); Zhou et al. (2023), computer vision (CV) Roser (2022); Haenlein and Kaplan (2019), natural sciences Han et al. (2023); Ivanenkov et al. (2023), etc. With the advent of neural networks, which are, unfortunately, uninterpretable black boxes Alwosheel et al. (2021), the scientific and industrial communities began to pay attention to the issues of stability and robustness Hamon et al. (2020); Kejriwal et al. (2024). First, let us define what we call ’robustness’ hereafter. It is known, that negligible changes in the input data (both for training model or inference model) can provoke uncontrolled changes in the output of AI. Since the AI-assisted solutions affect real world decisions, this output, in turn, can lead to erroneous, costly, unethical, or even dangerous results. Thus, the robustness of machine learning (ML) algorithms characterize absence of the aforementioned situations and seems to be one of the key issues in the design of AI-assisted systems. The similar problems were discussed in detail by researchers in the NLP Goyal et al. (2023); Wang et al. (2021) and CV Liu and Jin (2023); Li et al. (2024) domains, but it can certainly arise wherever neural networks are used, including materials science.
@@
Indeed, AI fever has not spared materials science either: ML has become a solid part of the pipeline for searching for new materials by high-throughput thermodynamic stability assessments within the vast chemical search spaces. The most recent approaches in the field are resting upon testing chemical modifications of existing materials through their compositional/chemical modifications Merchant et al. (2023); Chen et al. (2024) instead of generation of crystal structures from scratch. Despite the fact that many modern computational datasets, such as that of Materials Project Jain et al. (2013) and AFLOWLib Curtarolo et al. (2012a), can be used to pre-train AI models for their subsequent applications, fine-tuning stage or even more sophisticated routines (e.g., active learning Merchant et al. (2023); Yuan et al. (2023)) are generally required for disordered materials. In such approaches, the more complex crystal structure, the more complex the target search space. For defects in 2D materials modelled within the supercell approach Huang et al. (2023), the authors provided combinations of structured composition/configuration spaces at low defect contents and random structures subsamples at higher ones. Nevertheless, it was recently demonstrated Eremin et al. (2024) that physicochemical intuition can provide an acceptable quality of model predictions at low computational costs. Thus, training data sampling not only can play a major role in most effectively combining exhaustive and high-precision (such as density functional theory) calculations and fast screening data-driven approaches, but is also a point for potential instability: small errors and inaccuracies in the composition of the training dataset can lead to unnecessary and unsuccessful resource-intensive ab-initio computational and expensive real-life experiments.
@@
In our recent research Eremin et al. (2022), we addressed thermodynamic stability assessments for the disordered Sc-rich complex (ca. 140 atoms/cell) intermetallics by combining density functional theory calculations and data-driven approaches. During the data collection, there were obtained considerable differences in thermodynamic favorability of defects of certain types. Despite this fact, all the data collected was used for model training. This allowed graph neural networks (GNN) and descriptor-based ML models to gain sufficient generalization ability. In turn, at the inference stage, the complete composition/configuration space was built only for the thermodynamically favored defects resulting in an appropriate complexity of target search space. Thus, the peculiarity of the developed approach was that structures with disorder of different types were used at the training/validation and inference stages.
@@
The aim of the current contribution is to estimate the instability of a screening pipeline and evaluate robustness of different backbones of it. The main goal is evaluating impact of the less relevant to the target domain data impact on the results. Using the data and the approach presented in Eremin et al. (2022), we have generated 13 subdatasets based on the disordering level and thus have modeled the changes in the training data. It is shown that the Random Forest algorithm Breiman (2001) is robust and stable against the training data changes, while the Allegro Musaelian et al. (2023) neural network is not. Its predictions could become completely different; the assessment was done not only in a quantitative root squared error (RMSE) way, but as well in a qualitative manner. We have also demonstrated that the pre-training on the corresponding Aflow dataset Curtarolo et al. (2012a) slice increases stability and introduced an elementary sequential training trick to increase algorithm robustness without involving additional data.
@@ 3.1 Overall description
The dataset consists of doped Sc-M intermetallic systems, scandium-rich Mackay-type quasicrystal approximants. The dopants M are four metals: Pt, Pd, Ir and Rh, and thus four corresponding subdatasets were formed. On the A4 and A5 sites corresponding to the intercluster regions, the original Sc60M13 crystal structure was disordered by adding vacancies or Sc substitutions by metal M with a step of 2 atoms. Then the structures with low Sc content (less than 0.667 for Pt and Pd, less than 0.733 for Ir and Rh) were excluded from the further consideration due to the well-known competing phases. For the rest part of the search space, one random realization of crystal structure with each of the remaining compositions was chosen and computed using Vienna Ab initio Simulation Package (VASP) Kresse and Hafner (1993) Density Function Theory (DFT)-solver.
@@ 3.2 The term "limit"
The DFT-calculations showed that defects on the A4 site are more energetically favorable than that on A5 one, and thus we used numbers of defects on A5 as a quantitative measurement of sample similarity. So, we introduce the ’limit’ term corresponding to the maximum number of defects on the A5 site allowed to add to a certain subdataset. The maximum limit in the dataset equals the multiplicity of the A5 site, 24, and the minimum is 0. Thus, the entire dataset can be represented as 13 nested datasets, with the limit from 0 to 24. The obtained subdataset sizes depending on the limit are shown in Figure 1.
@@ 3.3 Splits
Secondly, each of the four Sc-Pd, Sc-Pt, Sc-Ir, and Sc-Rh datasets is divided into train, validation, and test samples in the same manner: the test set consists only of limit=0 part, while train and val contain all the other systems with the number of defects on A5 from 0 to 24. Train/val partition was made once for the full (limit=24) dataset, and then applied accordingly to the subdatasets with limits less than 24. The precise sizes of all the data samples obtained are represented in Table 1
@@ 3.4 Target energies
It is worth noting that the energy above the hull is, by design, positive for most systems, and that the ultimate goal of the screening pipeline is to search for crystals that have negative (or at least slightly positive) energy: only such systems meets synthesizability restrictions. Figure 2 illustrates the distribution of the energies above the hull in the Sc-Pd combined dataset.
Figure 2: Histogram of energies above the hull histogram for the Sc-Pd dataset
@@ 4 Methods
For each architecture and for each dataset, we used the following procedure: 13 models were trained (one for each limit=$x$ subdataset) and compared on the single test sample. Then an additional comparison was made on the inference dataset
@@ 4.3 Training
For classical models, 5-fold cross-validation was used the predictions of which were then averaged, as for the neural network we used a described in 3.3 piece of data for validation purposes. Root-mean-square error (RMSE) 3 was used as error function for training and validation steps. Random seeds were fixed for all experiments.
@@ 5.1 Test
Even though RMSEs of all classical Gradient Boosting runs (see Figure 3) does not significantly vary (XGBoost) or have a clearly visible pattern (CatBoost and LightGBM),
@@
all the chosen Gradient Boosting algorithms perform worse than Random Forest approach, which can be seen from Figure 4
@@
Random Forest have much lower test RMSE compared to that of the tested Gradient Boosting models. Random Forest shows test RMSE of ca. 0.01 eV/atom, and Gradient Boosting models’ scores are of ca 0.02 eV/atom. The latter have a tendency to overfit on the subdatasets with bigger limits (while test sample corresponds to limit=0) and their predictions are highly biased. Moreover, we checked made and checked the predictions of the trained Gradient Boosting models for the inference set. The results obtained were rather unrealistic and inconsistent. For this reason, Gradient Boostings were excluded from further consideration and only the Random Forest model is used.
Random Forest provides stable results, i.e., the test RMSEs are found to fluctuate within a few meV/atom for three out of four datasets (except for the Sc-Pd systems) as shown in Figure 5.
@@
In Figure 6, the SimilarityScores matrices also illustrate the Random Forest stability: sharp changes are common only for the Sc-Pt systems.
@@
Different neural network training approaches perform about the same RMSE ranges of 0.010–0.015 eV/atom on the test set. As it is shown in Figure 7, the self-pre-trained Allegro model’s errors are always smaller than that of the ordinarily trained one. The Aflow-pretrained version, in turn, shows bigger errors.
@@
In contrast to the similar RMSE values of the models tested, the Similarity Score shows pronounce differences between the model predictions at a qualitative level. The Aflow- (Figure 8) and self-pretrained (Figure 9) Allegro models provide heatmaps that illustrates larger Similarity Scores than that on the ordinarily trained network (Figure 10).
@@
Moreover, for the self-pretrained Allegro the light areas are much more clearly located. If a light bar appears after a certain limit, it means that it is likely to continue at the next limits. For some metals, there is a pattern where the model gets good results up to a certain limit, comprising the relevant data only. Then, the model shows a worse result due to getting a little less relevant data. However, the model being provided by a large amount of both the relevant (limit=0) and less relevant (limit=$x$, $x>0$) data shows better results again. It can be assumed that by perceiving the overall data, it increases the robustness of predictions.
@@ 5.2 Inference
The comparison between the limit=24 models and all the others shows that the Random Forest is robust, and the Allegro model is not. The recall metric computed considering the limit=24 predictions as the ground truth is always 1 for Random Forests (see Figure 11 a) and vary from 0 to 1 for Allegro (see Figure 11 b).
@@
The pairwise comparison of limits described in 4.5 highlights the apparent difference between the models, trained on different subdatasets with limit=$x$. Random Forest behaves stable, showing not the chaotic changes, but a visible transition (see Figure 12) from different to similar runs. The empty graphs and solid-colored heatmaps in Figure 12 show that there are no structures predicted below the convex hull.
@@
The heatmaps indicate the instability of neural networks, as this behavior is not observed in the results of the Random Forest. Despite the fact that the results may vary, they are not random, and there is some kind of dependence factor. As it can be seen in Figure 13, there is a trade-off between the relevance of certain data points and their amounts provided to the model. In the middle range of the introduced limits, a small change in the limit leads to large changes in the results. This observation clearly illustrates the following principle – training data should be either small and strictly relevant or big and less relevant. From the dependencies obtained for Random Forest, this can also be seen.
@@
As shown in Figures 14, 15, the trade-off is definitely more visible and after a certain limit on each metal there is a great similarity between structures that are predicted below the convex hull. This also shows that there is a certain “relevance threshold” for palladium and platinum. With pretrained network, we can overcome unstable behavior of neural-network-based approaches without generation of additional data. Thresholds differ from those seen in the pictures of sequential training and regular training. This is most likely due to the fact that the model was already familiar with similar data patterns, and its ordinary training requires less data particular for a certain task. For rhodium, the threshold is also apparently observed, which shows that when training on data with a limit of 6 or more, the model predicts structures below the hull. The results for iridium do not allow us to unambiguously assess the effectiveness of the method proposed.